Consider an RC series circuit with e=24volts, R=10, and C=15.calculate the following (I) the maximum charge on the capacitor, (ii) the time constant of the circuit (iii) the initial current if the internal resistance, r of the battery is 2
1
Expert's answer
2018-08-13T12:08:08-0400
I ) q=C·U_c=15·〖10〗^(-6)·24=0.36·〖10〗^(-3) C

ii ) τ=R·C=10·15·〖10〗^(-6) 0.15 ms

iii) I_initial=e/(R+r)=24/(10+2)=2 A